OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 3(2x - 1) greater than or equal to 4(2x - 3) - 3? x less than or equal to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

6*x - 3 >= 8*x - 15

OpenStudy (anonymous):

12 >= 2*x

OpenStudy (anonymous):

6 >= x

OpenStudy (anonymous):

No, the last choice is correct

OpenStudy (anonymous):

Tell us why...

OpenStudy (anonymous):

it would be c

OpenStudy (anonymous):

wait how is it c?

OpenStudy (anonymous):

dont give the girl wron answers @cblrtopas its annoying

OpenStudy (anonymous):

ok hold on lol

OpenStudy (anonymous):

It's not wrong

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

could you both explain how you got the answer?

OpenStudy (anonymous):

I simplified both sides of the inequality until I had x on one side and numbers on the other

OpenStudy (jdoe0001):

well, your inequality looks like $$ 3(2x - 1) \ge 4(2x - 3) - 3 \implies 6x-3 \ge 8x-15 $$

OpenStudy (anonymous):

The only answer that is not the right way to represent the solution is shaded circle at 6 with shading to the right...

OpenStudy (jdoe0001):

now if you were to solve -> 6x-3 = 8x - 15 for "x" what would you get?

OpenStudy (anonymous):

I.e. the last response

OpenStudy (anonymous):

i woud get -2x=-12?

OpenStudy (jdoe0001):

right, and x = 6, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x is less than or equal to 6

OpenStudy (jdoe0001):

ok, in an inequality is the same procedure, with the only difference that, when you multiply or divide or exponentialize by a negative number, like you just did now, you have to \(\color{red}{flip}\) the sign

OpenStudy (anonymous):

yes i remember that u have to flip the sign

OpenStudy (jdoe0001):

$$ 3(2x - 1) \ge 4(2x - 3) - 3 \implies 6x-3 \ge 8x-15\\ -2x \ge -12 \implies x \le 6 $$

OpenStudy (anonymous):

so the answer would be the last choice?

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):



OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

see the solid line? see the shading?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so, that's the \(x \le 6\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so wait im confused whats the answer

OpenStudy (jdoe0001):

well, let's see the possible choices $$ \text{ A number line with a closed circle on 6 and shading to the} \color{red}{\text{ left}}\\ \\ \text{ A number line with a closed circle on 6 and shading to the} \color{red}{\text{ right}} $$ so, which one do you think?

OpenStudy (anonymous):

the second one?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

its to the left but gl with that

OpenStudy (anonymous):

ok the solution to the inequality is: x are all numbers less than or equal to the number 6.... Your question asks for what is the opposite of the solution to the inequality...essentially. The opposite of x <= 6 is of course x >= 6 and the statement: "A number line with a closed circle at 6 and shading to the right."